# MIXRTs: Toward Interpretable Multi-Agent Reinforcement Learning via Mixing Recurrent Soft Decision Trees

### v3

@@ 3 Recurrent Tree Cells
In this section, we propose RTCs that introduces recurrency into SDTs to encode the previous information and enhance the fidelity of the Q-value with the ensemble technique. First, we propose an RTC that receives the current individual observation and relies on historical information to capture long-term dependencies in partially observable tasks. Then, we utilize a linear combination of multiple RTCs via an ensemble mechanism to yield high performance and reduce the variance of the model while retaining simplicity and interpretability.
@@ 3.1 Recurrent Tree Cell
The complexity of neural networks brings barriers to understanding, verifying, and trusting the behavior of the agent, as it has complex transformations and nests non-linear operators. To relieve the dilemma, SDT offers an effective way to interpret the decision-making pathways by visualizing decision nodes and associated probabilities. As depicted in Fig. 1(a), SDT is an univariate differentiable tree with a probabilistic decision boundary at each filter. However, SDT has no explicit mechanisms for deciphering the underlying state of POMDP in sequential decision-making tasks over longer timescales. Generally, it could degrade the performance by estimating the Q-value from an incomplete observation instead of a global state. Indeed, by leveraging recurrency to a deep Q-network with a recurrent neural network (RNN), agents can effectively capture long-term dependencies conditioning on their entire action-observation history [47]. Inspired by the advances in the recurrent architecture, we first introduce recurrency into the SDT structure and propose a recurrent tree cell (RTC) to encode long-term dependencies. As shown in Fig. 1(b), RTC receives the current individual observation $o^{t}_{i}$ and history embedding $h^{t-1}_{i}$ as input at each time step, and extracts the information of the previous hidden state for each agent $i$.
Combining the recurrent architecture, for each non-leaf node, RTC learns linear filters in its branching node to traverse to the left child node with the probability as
@@
where $b^{j}$ is a learnable bias parameter, and $w^{j}_{o}$ and $w^{j}_{h}$ are the learnable parameters for the feature vectors of the observation $o^{t}_{i}$ and the previous hidden state $h^{t-1}_{i}$, respectively. Similarly, we can obtain the probability of the right branch as $1-p_{j}(o_{i}^{t},h_{i}^{t-1})$, which is the same as in SDT. Then, for each leaf node, the probability of selecting the leaf $P^{l}$ is equal to the product of the overall path probability starting from the root to the leaf node $l$ as
$P^{l}(o_{i}^{t},h_{i}^{t-1})=\prod_{j\in route(l)}\left\{p_{\left\lfloor j/2\right\rfloor\rightarrow j}(o_{i}^{t},h_{i}^{t-1})^{\left[l\swarrow j\right]}\,\right.$ (7)
$\left.(1-p_{\left\lfloor j/2\right\rfloor\rightarrow j}(o_{i}^{t},h_{i}^{t-1}))^{1-\left[l\swarrow j\right]}\right\},$
where $\left[l\swarrow j\right]$ denotes the boolean event $\in\left\{0,1\right\}$, which means whether $j$ passes the left-subtree to reach the leaf node $l$. With the target distribution of the tree, we measure the current hidden state $h_{i}^{t}$ of the leaf by combining the probability values of each leaf with the weight scalar $\theta^{l}_{h}$ and provide a vector $w_{q}$ that serves this tree to capture the action-observation value as
@@
$Q_{i}(\tau_{i},\cdot)=$ $~{}~{}~{}~{}~{}~{}h_{i}^{t}\,w_{q},$
where $P^{l}$ is the overall path probability along the root to the leaf $l$. Different from the leaf nodes of SDT, the learnable parameter $\theta^{l}_{h}\in\mathbb{R}$ in an RTC calculates the hidden state $h_{i}^{t}$ by multiplying with $P^{l}$, and $w_{q}\in\mathbb{R}^{\mathcal{U}}$ is a training parameter to transform the hidden information $h_{i}^{t}$ into the action distribution. Finally, in a decentralized execution way, we choose an action $u_{i}^{t}$ for each agent $i$ by $\varepsilon$-greedy methods with respect to its $Q_{i}(\tau_{i},u_{i})$. This decision process retains the simplicity and interpretability of the model since it can present which dimensions of observation influence the action distribution during the inference.
@@ 3.2 Ensemble of Recurrent Tree Cells
The SDT-based methods exhibit several appealing properties with the multivariate tree structure, such as ease of tuning, robustness to outliers and good interpretability [31]. Since all input features are used for each node with a multivariate setting, a single SDT could lack expressiveness and output predictions with high variance. An ensemble paradigm of models, rather than a single one, is well known to reduce the variance component of the estimation error [48, 49]. To relieve the above tension while maintaining the interpretability, we linearly combine multiple RTCs with the variance-optimized bagging (Vogging) approach based on the bagging technique [50]. Based on the Vogging ensemble mechanism, the individual value function for the agent $i$ can be represented as
$Q_{i}(\tau_{i},\cdot)=h_{i,(1)}^{t}\,w_{q,(1)}+h_{i,(2)}^{t}\,w_{q,(2)}+...+h_{i,(H)}^{t}\,w_{q,(H)},$ (9)
where $H$ is the size of ensemble RTCs, $w_{q,(1)},w_{q,(2)},...,w_{q,(H)}$ are learnable parameters to optimize a linear combination of the $H$ trees for improving the expressiveness and reducing the prediction variance of RTCs, and the hidden state $h^{t}_{i}\in\mathbb{R}^{H}$ is derived as the vector $[h_{i,(1)}^{t},h_{i,(2)}^{t},...,h_{i,(H)}^{t}]$ to express the history record. Finally, the RTCs produces a local policy distribution as a Q-function for each agent and execute the action via sampling the distribution. Here, we utilize back-propagating gradients to update the $Q_{i}$ from the Q-learning rule [36, 39] via the joint reward over all agents.
@@
The functions used for all sub-modules in the RTCs ensemble are linear, which both maintains the interpretability of the model and simplifies it. Note that when the model structure gets complex, the interpretability decreases as the simplicity of the model is lost. Since shallow trees can be easily described and even be implemented by hand, it is acceptable that linearly combining the trees of depth $2$ or $3$ achieves acceptable performance and stability via only sacrificing a little bit model interpretability. In practice, a moderate number of ensemble trees $H$ (e.g., $16\sim 64$) is sufficient to obtain efficient performance, and additional ablation results are provided in Section 8.
@@ 4 Mixing Recurrent SDTs Architecture
Figure 2: MIXRTs architecture. (a) Diagram of the structure of the mixing tree with depth $2$. (b) In the overall MIXRTs architecture, we finally obtain the joint $Q_{tot}$ value via a linear combination of the individual action-value functions. (c) An individual RTCs structure for each agent.
In this section, we propose a novel method called MIXing Recurrent soft decision Trees (MIXRTs), which can represent a much richer class of action-value functions analogous to the advanced CTDE algorithms. The overall architecture of MIXRTs is illustrated in Fig. 2, including two main components as follows: (i) an ensemble of RTCs as an individual action-value function for each agent, and (ii) a mixing component similar to the ensemble of RTCs in which a joint action-value function $Q_{tot}$ is factorized into the individual action-value function $Q_{i}$ of each agent $i$ under the IGM constraint. The two components are trained in a centralized manner and each agent uses its own factorized individual action-value function $Q_{i}$ to take actions during decentralized execution. Each component is elaborated on next.
@@ 4.1 Individual RTCs for the Action Value
For each agent, the individual action-value function generally could be represented by an ensemble of RTCs, where different agents are mutually independent RTCs. However, the joint action-observation space grows exponentially as the number of agents increases, and then its large number of learnable parameters may cause confusion in understanding the decision process of the model. Inspired by parameters sharing [2], we utilize the ensemble of RTCs to share certain parameters of RTCs between agents for improving the learning efficiency, which can also give rise to the concept of agent invariance and help avoid the lazy agent problem. To keep the agent invariance, we incorporate the role information into the agent via a one-hot encoding of its identity, and the role information is concatenated with the corresponding observation at the root layer. The architecture, with shared weights and information channels across agent networks, satisfies agent invariance with the identical policy and improves simplicity and interpretability with fewer parameters. For each agent $i$, the individual value is represented by an ensemble of RTCs, which takes the current local observation $o^{t}_{i}$, the previous hidden state $h^{t-1}_{i}$ and the previous action $u^{t-1}_{i}$ as the inputs, and then outputs the local action-value $Q_{i}(\tau_{i},u_{i})$.
@@ 4.2 Mixing Trees Architecture
The mixing trees architecture follows an interpretable Q-value decomposition formation that connects the local and joint action-value functions via feed-forward ensemble trees, which is also similar to the ensemble of RTCs without embedding the history information. It incorporates the information of global state $s_{t}$ into individual action-value functions during the centralized training process and produces the weights of the local action-value in the joint action-value $Q_{tot}$. The main assumption is that the joint action-value function can be additively decomposed into value functions across agents. However, unlike VDN or QMIX, we assume that the joint action-value function can be approximately weighted by the individual Q-values in a linear manner as
@@
where $\mathbf{\bm{\tau}}$ and $\bm{u}$ are joint action-observation and joint action, respectively, and we force the assignment of credit weights $W_{i}>0$. The choice of linear weighting aims to render a reliable interpretation of the relationship between the agents. Generally, the agent who receives a higher credit assignment with a larger weight $W_{i}$ makes a greater contribution to the team.
In detail, we employ the interpretable RTCs to obtain the weights $W_{i}$ for each agent $i$, which establishes the relations from the individuals to the global. Each root of the ensemble of RTCs takes the individual action-value $Q_{i}$ and the global state vector $s^{t}$ as input as
@@
where $w^{j}_{q}$, $w^{j}_{s}$, and $b^{j}$ are learnable parameters for each layer of the RTC. After the inference process, we can obtain the joint probability distributions $P^{l}$ at the leaf node layer. Here, we utilize the same operation as the individual action-value RTCs with $H$ trees to obtain the stable weights as
@@
$W_{i}=$ $\frac{exp(\sum_{k=1}^{H}\phi_{i,(k)}\,w_{\phi,(k)})}{\sum_{i=1}^{n}exp(\sum_{k=1}^{H}\phi_{i,(k)}\,w_{\phi,(k)})},$
where $\theta^{l}_{i},w_{\phi}$ are the learnable parameters and the mixing weight $W_{i}$ is positive with softmax operation to enforce the monotonicity between $Q_{tot}$ and each $Q_{i}$. Here, the ultimate goal of the mixing architecture is not only to describe the realization of efficient value decomposition, but also to elaborate an intuitive interpretation that produces the importance of observation attributes and the weight of each agent at different time steps for one episode. Further, Theorem 1 presents the theoretical analysis that MIXRTs guarantees the monotonicity by imposing an easy constraint on the relationship between $Q_{tot}$ and each $Q_{i}$, which is sufficient to satisfy Eq. 4. This allows each agent $i$ to participate in a decentralized execution solely by choosing greedy actions.
@@
All parameters $\mathit{\Theta}$ in MIXRTs are learned by sampling a number of transitions $b$ from the buffer $\mathcal{B}$ and minimizing the following expected squared temporal difference (TD) error loss as
@@
and the target is obtained via Double DQN [51, 52], which estimates as $y^{\prime}=r+\gamma\overline{Q}_{tot}(\bm{\tau^{\prime}},\arg\max_{\bm{u^{\prime}}\in\mathcal{U}^{n}}{Q}_{tot}(\mathbf{\bm{\tau^{\prime}}},\bm{u^{\prime}};\mathit{\Theta^{\prime}}))$. $\mathit{\Theta^{\prime}}$ are the parameters of a target network that are periodically copied from $\mathit{\Theta}$. Since Eq. (4) holds, we maximize $Q_{tot}$ in a linear fashion to achieve competitive performance, rather than scaling the non-linear case.
@@
2s3z 2 Stalkers, 3 Zealots 2 Stalkers, 3 Zealots $1500M$
@@ 5 Experiments
In this section, we evaluate MIXRTs on two representative benchmarks as our testbed: the Multi-agent Particle Environment (MPE) [53] and StarCraft Multi-Agent Challenge (SMAC) [33]. The goal of our experiments is to evaluate the performance and demonstrate the interpretability of MIXRTs. We compare our method with the widely investigated algorithms, including VDN [2], QMIX [3], QTRAN [37] and QPLEX [34], since they are advanced value-based methods that train decentralized policies in a centralized fashion. Note that our focus is to strike a balance between model interpretability and learning performance, other than blindly beating state-of-the-art baselines. Additionally, we compare with existing interpretable models SDTs [27] and CDTs [31] that decompose the multi-agent task into a set of simultaneous single-agent ones [44, 54]. Further, we perform the sensitivity analysis on two key hyperparameters: the tree depth and the number of trees. Finally, we concretely present the interpretability of MIXRTs in regard to the learned tree model at SMAC, aiming to present an easy-to-understand decision-making process from the captured knowledge of the tasks.

### v4

@@ 3 Recurrent Tree Cells
In this section, we propose RTCs, which introduce recurrency into SDTs to encode historical information and enhance the fidelity of the Q-value through the ensemble technique. First, we propose an RTC that receives the current individual observation and relies on historical information to capture long-term dependencies in partially observable tasks. Then, we utilize a linear combination of multiple RTCs via an ensemble framework to yield high performance and reduce the variance of the model, all while retaining simplicity and interpretability.
@@ 3.1 Recurrent Tree Cell
The complexity of neural networks brings barriers to understanding, verifying, and trusting the behaviors of agents, as it has complex transformations and non-linear operators. To relieve this dilemma, the SDT offers an effective way to interpret the decision-making pathways by visualizing decision nodes and associated probabilities. As depicted in Fig. 1(a), an SDT shows an univariate differentiable tree with a probabilistic decision boundary at each filter. However, the SDT lacks explicit mechanisms for deciphering the underlying state of POMDP in sequential decision-making tasks over longer timescales. This limitation can degrade performance by estimating the Q-value from incomplete observations rather than the global state. Indeed, by leveraging recurrency to a deep Q-network [47], a recurrent neural network (RNN) can effectively capture long-term dependencies conditioning on their entire action-observation history. Inspired by advanced recurrent architectures, we first introduce recurrence into an SDT and propose a recurrent tree cell (RTC) to capture long-term dependencies. As shown in Fig. 1(b), an RTC receives the current individual observation $o^{t}_{i}$ and the previous embedding $h^{t-1}_{i}$ as input at each time step, extracting hidden state information for each agent $i$.
For an RTC, each non-leaf node learns a linear filter to traverse to its left child node with the probability as
@@
where $w^{j}_{o}$ and $w^{j}_{h}$ are learnable parameters for the observation $o^{t}_{i}$ and the previous hidden state $h^{t-1}_{i}$, respectively, and $b^{j}$ is a learnable bias. Similarly, the probability of traversing to the right child is $1-p_{j}(o_{i}^{t},h_{i}^{t-1})$.
We now discuss how to model the probability that an observation reaches a selected leaf $l$. Let $\left[l\swarrow j\right]$ denotes a boolean event $\in\left\{0,1\right\}$, indicating whether $j$ passes the left-subtree to reach the leaf node $l$. The probability $P^{l}$ that the observation $o_{i}^{t}$ reaches $l$ for each agent $i$ is given by
$P^{l}(o_{i}^{t},h_{i}^{t-1})=\prod_{j\in route(l)}\left\{p_{j}(o_{i}^{t},h_{i}^{t-1})^{\left[l\swarrow j\right]}\,\right.$ (4)
$\left.(1-p_{j}(o_{i}^{t},h_{i}^{t-1}))^{1-\left[l\swarrow j\right]}\right\},$
where $p_{j}$ is the probability of going from a current node $j$ to its left child node $2j$, as illustrated in Fig. 1.
With the target distribution of the tree, we measure the current hidden state $h_{i}^{t}$ of the leaf by combining the probability values of each leaf with a scalar weight $\theta^{l}_{h}$, and provide a vector $w_{q}$ that serves this tree to capture the action-observation value as
@@
$Q_{i}(\tau_{i},\cdot)=$ $~{}~{}~{}~{}~{}~{}w_{q}^{\top}h_{i}^{t},$
where $P^{l}$ is the overall path probability along the root to leaf $l$. Different from the leaf nodes of the SDT, each learnable parameter $\theta^{l}_{h}\in\mathbb{R}$ calculates the hidden state $h_{i}^{t}$ by weighting with $P^{l}$, and $w_{q}\in\mathbb{R}^{1\times\mathcal{U}}$ is a training parameter to transform the hidden state $h_{i}^{t}$ into the action distribution in an RTC. Finally, during decentralized execution, each agent $i$ selects an action $u_{i}^{t}$ using $\varepsilon$-greedy strategy with respect to its estimated $Q_{i}(\tau_{i},u_{i})$. This decision process retains the simplicity and interpretability of the model, as it reveals which dimensions of the observation influence the action distribution during the inference.
@@ 3.2 Ensemble of Recurrent Tree Cells
SDT-based methods exhibit several appealing properties in multivariate tree structures, such as ease of tuning, robustness to outliers, and good interpretability [32]. Since all input features are used for each node in a multivariate setting, a single SDT may lack expressiveness and output predictions with high variance. It is well known that an ensemble of models can reduce the variance component of the estimation error [48, 49]. To relieve the above tension while maintaining the interpretability, we linearly combine multiple RTCs with the variance-optimized bagging (Vogging) approach based on the bagging technique [50]. Based on the Vogging ensemble mechanism, the individual value function for the agent $i$ can be represented as
$Q_{i}(\tau_{i},\cdot)=w_{q,(1)}^{\top}\,h_{i,(1)}^{t}+w_{q,(2)}^{\top}\,h_{i,(2)}^{t}+...+w_{q,(H)}^{\top}\,h_{i,(H)}^{t},$ (6)
where $H$ is the size of ensemble RTCs, $w_{q,(1)},w_{q,(2)},...,w_{q,(H)}$ are learnable parameters used to optimize the linear combination of the $H$ trees for improving the expressiveness and reducing the prediction variance of RTCs, and the hidden state $h^{t}_{i}\in\mathbb{R}^{H\times 1}$ is rewritten as the vector $[h_{i,(1)}^{t},h_{i,(2)}^{t},...,h_{i,(H)}^{t}]^{\top}$ to represent the history record, i.e., $h_{i}$ is a representation of the history $\tau_{i}$. Finally, RTCs produce a local policy distribution as a Q-function for each agent and execute an action by sampling this distribution. Here, the loss function of RTCs working on a single agent is based on Q-learning [37, 40], updating $Q_{i}$ as
$\mathcal{L}=\mathbb{E}\left[\left(r_{i}+\gamma\max_{u^{\prime}}Q_{i}(\tau_{i}^{\prime},u^{\prime})-Q_{i}(\tau_{i},u)\right)^{2}\right],$ (7)
where $r_{i}$ is the reward received by agent $i$. However, in multi-agent systems, all agents only share a team reward $r$, and therefore we explore the value decomposition in Section 4.
@@
The functions used for all sub-modules in the RTCs are linear, which both preserves the interpretability of the model and simplifies its structure. It is important to note that as the model becomes more complex, interpretability diminishes due to the loss of simplicity. Since shallow trees can be easily described and even be implemented manually, it is acceptable that linearly combining trees of depth $2$ or $3$ achieves acceptable performance and stability while only sacrificing a minimal degree of interpretability. In practice, a moderate number of ensemble trees $H$ (e.g., ranging from $16\sim 64$) is sufficient to obtain efficient performance, and additional ablation results are provided in Section 8.
@@ 4 Mixing Recurrent SDTs Architecture
Figure 2: MIXRTs architecture. (a) Diagram of the structure of the mixing tree with depth $2$. (b) In the overall MIXRTs architecture, we finally obtain the joint $Q_{tot}$ value via a linear combination of the individual action-value functions. (c) Individual RTCs for each agent.
In this section, we propose a novel method called MIXing Recurrent soft decision Trees (MIXRTs), which can represent a much richer class of action-value functions analogous to advanced CTDE algorithms. The overall architecture of MIXRTs is illustrated in Fig. 2, including two main components as follows: (i) an ensemble of RTCs as an individual action-value function for each agent, and (ii) a mixing component similar to the ensemble of RTCs in which a joint action-value function $Q_{tot}$ is factorized into the individual action-value function $Q_{i}$ of each agent $i$ under the IGM constraint. The two components are trained in a centralized manner and each agent uses its own factorized individual action-value function $Q_{i}$ to take actions during decentralized execution. Each component is elaborated on next.
@@ 4.1 Individual RTCs for the Action Value
For each agent, the individual action-value function can generally be represented by an ensemble of RTCs, where different agents are mutually independent RTCs. However, as the number of agents increases, the joint action-observation space grows exponentially, leading to a large number of learnable parameters that may confuse understanding the decision process of the model. Inspired by parameters sharing [2], we utilize the ensemble of RTCs with shared parameters among agents to improve the learning efficiency. This operation can also give rise to the concept of agent invariance and help avoid the lazy agent problem. To keep agent invariance, we incorporate role information into each agent via a one-hot encoding of its identity, which is concatenated with the corresponding observation at the root layer. The architecture, with shared weights and information channels across agent networks, satisfies agent invariance with the identical policy, thereby enhancing simplicity and interpretability with fewer parameters. For each agent $i$, the individual value is represented by an ensemble of RTCs that takes the current local observation $o^{t}_{i}$, the previous hidden state $h^{t-1}_{i}$ and the previous action $u^{t-1}_{i}$ as inputs, and then outputs the individual action value $Q_{i}(\tau_{i},u_{i})$.
@@ 4.2 Mixing Trees Architecture
The mixing trees architecture follows an interpretable Q-value decomposition that connects the local and joint action-value functions via feed-forward ensemble trees, which is similar to the ensemble of RTCs but without embedding the history information. During the centralized training process, it incorporates the information of global state $s_{t}$ into individual action-value functions and produces the weights of the local action-value in the joint action-value $Q_{tot}$. The main assumption is that the joint action-value function can be additively decomposed into value functions across agents. However, unlike VDN or QMIX, we assume that the joint action-value function can be approximated as a linear weighting of the individual Q-values
@@
where $\mathbf{\bm{\tau}}$ and $\bm{u}$ are joint action-observation and joint action, respectively, and we force the assignment of positive credit weights $W_{i}>0$. The linear weighting is chosen to provide a reliable interpretation of the relationships among agents. In general, the agent assigned a higher credit assignment with a larger weight $W_{i}$ makes a greater contribution to the team.
In detail, we employ the interpretable RTCs to obtain weights $W_{i}$ for each agent $i$, which establish relations from the individuals to the global. Each filter of RTCs takes the individual action value $Q_{i}$ and the global state $s^{t}$ as inputs, and calculates the probability of transitioning to the left child node as
@@
where $w^{j}_{q}$, $w^{j}_{s}$, and $b^{j}$ are learnable parameters for each node of RTCs. After the inference process, we obtain the joint probability distributions $P^{l}$ at the leaf layer. To achieve stable weights, we employ an ensemble mechanism with $H$ trees, resulting in
@@
$W_{i}=$ $\frac{\exp(\sum_{k=1}^{H}w_{\phi,(k)}^{\top}\,\phi_{i,(k)})}{\sum_{i=1}^{n}\exp(\sum_{k=1}^{H}w_{\phi,(k)}^{\top}\,\phi_{i,(k)})},$
where $\theta^{l}_{i}\in\mathbb{R}$ and $w_{\phi}\in\mathbb{R}^{1\times n}$ are learnable parameters in one tree. The mixing weight $W_{i}$ is ensured to be positive through the use of softmax operation, which enforces monotonicity between $Q_{tot}$ and each $Q_{i}$. Here, the goal of the mixing architecture is not only to achieve efficient value decomposition but also to provide an intuitive interpretation that highlights the importance of observation attributes and the weights of each agent at different time steps within an episode. Further, Theorem 1 provides a theoretical analysis showing that MIXRTs guarantee monotonicity by imposing a simple constraint on the relationship between $Q_{tot}$ and each $Q_{i}$, which is sufficient to satisfy Eq. 1. This allows each agent $i$ to participate in a decentralized execution solely by choosing greedy actions.
@@
All parameters $\mathit{\Theta}$ in MIXRTs are learned by sampling a batch of transitions $b$ from the buffer $\mathcal{B}$ and minimizing the following expected squared TD error loss as
@@
where the target value $y^{\prime}$ is obtained using Double DQN [51, 52] and is estimated as $y^{\prime}=r+\gamma\overline{Q}_{tot}(\bm{\tau^{\prime}},\arg\max_{\bm{u^{\prime}}\in\mathcal{U}^{n}}{Q}_{tot}(\mathbf{\bm{\tau^{\prime}}},\bm{u^{\prime}};\mathit{\Theta^{\prime}}))$. Here, $\mathit{\Theta^{\prime}}$ denotes all parameters of the target network, which are periodically copied from $\mathit{\Theta}$. Since Eq. (1) holds, we maximize $Q_{tot}$ in a linear fashion to achieve competitive performance without resorting to more complex non-linear methods.
@@
2s3z 2 Stalkers, 3 Zealots 2 Stalkers, 3 Zealots $1500K$
@@ 5 Experiments
In this section, we evaluate MIXRTs on two representative benchmarks as our testbed: the Multi-agent Particle Environment (MPE) [53] and StarCraft Multi-Agent Challenge (SMAC) [34]. The goal of our experiments is to evaluate the performance and demonstrate the interpretability of MIXRTs. We compare our method with widely investigated algorithms, including VDN [2], QMIX [3], QTRAN [38] and QPLEX [35], since these are advanced value-based methods that train decentralized policies in a centralized fashion. It is important to note that our focus is on balancing model interpretability and learning performance, rather than blindly beating state-of-the-art baselines. Additionally, we compare MIXRTs with existing interpretable models SDTs [28] and CDTs [32], which decompose a multi-agent task into a set of simultaneous single-agent tasks [44, 54]. Further, we perform the sensitivity analysis on two key hyperparameters: the tree depth and the number of trees. Finally, we concretely present the interpretability of MIXRTs regarding the learned tree model at SMAC, aiming to present an easy-to-understand decision-making process from the captured knowledge of tasks.